given: N=250 n=100 μ=35.6 σ=4.89

find: p( ≥36)


1
Expert's answer
2021-07-15T04:33:40-0400

When samples are drawn from finite population


"z=\dfrac{x-\mu}{\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}}"

"z=\dfrac{36-35.6}{\dfrac{4.89}{\sqrt{100}}\sqrt{\dfrac{250-100}{250-1}}}=1.053914"

"P(X\geq36)=1-P(X<36)"

"\approx1-P(Z\leq1.053914)"

"\approx0.145961"